# Determination of the density in the linear elastic wave equation

### v2

@@ 4. S-waves and proof of the main theorem
$+\nabla\log\sqrt{\rho}\cdot\nabla c_{S}^{2}+\nabla_{\beta}\log\sqrt{\rho}\left(2\frac{c_{P}^{2}-c_{S}^{2}}{c_{S}^{4}}\nabla_{\alpha}c_{S}^{2}+c_{S}^{-2}\nabla_{\alpha}c_{P}^{2}-\frac{12}{c_{P}^{2}-c_{S}^{2}}\nabla_{\alpha}c_{S}^{2}\right)\eta^{\alpha}\eta^{\beta}$
+terms that do not depend on $+\text{terms that do not depend on $\rho$}$
$=$ $:N_{\alpha\beta}\eta^{\alpha}\eta^{\beta}+C,$
where $C$ does not depend on $\rho$. Note that the second order tensor $N$ can be written in a coordinate-free way
@@
$\begin{split}&\mathrm{i}c_{S}^{-2}\left(2\mu\partial_{3}a_{1\alpha}+(\partial_{3}\mu)a_{1\alpha}-\mu g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau}a_{1\beta}+\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\beta\gamma}}{\partial\tau}c_{S}^{-2}\mu a_{1\alpha}-c_{S}^{-1}\partial_{3}c_{S}\mu a_{1\alpha}\right)c_{S}^{2}\eta^{\alpha}\\ =&(\langle N,\eta\otimes\eta\rangle_{g_{S}}+C)J^{-1/2}c_{S}^{1/2}\rho^{1/2}.\end{split}$
Assume that $\gamma$ is an arbitrary geodesic in $(\Omega,g_{S})$ connecting two points $a\in\partial\Omega$ to $b\in\partial\Omega$, so the length of $\gamma$ is smaller than $T$. Denote $D_{\tau}$ to be the covariant derivative along the geodesic. We consider the above equation along $\gamma$. If we denote
$a_{1\alpha}=J^{-1/2}c_{S}^{1/2}\rho^{-1/2}\xi_{\alpha},$
with a vector field $\xi=(\xi_{1},\xi_{2},0)$ along $\gamma$, the above equation can be written as
@@
along $\gamma$.
Then
$\langle\xi,\eta\rangle_{g_{S}}\Big{|}_{a}^{b}=c_{S}^{2}(\xi\cdot\eta)\Big{|}_{a}^{b}=-\frac{\mathrm{i}}{2}\int_{\gamma}(\langle N,\eta\otimes\eta\rangle_{g_{S}}+C)\mathrm{d}s.$
So we can recover the transverse ray transform of $N$
$\int_{\gamma}\langle N(\gamma(s)),\eta\otimes\eta\rangle_{g_{S}}\mathrm{d}s$
from boundary value of $\mathbf{a}_{1}$ once the boundary value of $\rho$ is determined. Here $\eta$ could be an arbitrary parallel vector field along $\gamma$ and orthogonal to $\gamma$ with respect the metric $g_{S}$ (also with respect to the Euclidean metric $g_{E}$). Thus the transverse ray transform of $N$ can be recovered from the Dirichlet-to-Neumann map $\Lambda_{\rho}$. We note here that clearly the transverse ray transform of $N$ is equal to that of the symmetrization of $N$.
Next we work under the Cartesian coordinates. Under the strictly convex foliation condition, by the invertibility of the transverse ray transform [33], we can recover the symmetrization of $N$,
Here $I$ is the identity matrix (the Euclidean metric under Cartesian coordinates) and $\otimes^{s}$ denotes the symmetrized product
$u\otimes^{s}v=\frac{1}{2}(u\otimes v+v\otimes u).$
We also refer the readers to [21] for the invertibility of the transverse ray transforms under certain curvature conditions. In conclusion, if $\Lambda_{\rho_{1}}=\Lambda_{\rho_{2}}$, we have $\mathrm{Sym}(N_{1})=\mathrm{Sym}(N_{2})$, which can be written as a second order differential equation for $\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}}$,
$\begin{split}&\Delta(\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}})I+2\nabla^{2}(\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}})+(|\nabla\log\sqrt{\rho_{1}}|^{2}-|\nabla\log\sqrt{\rho_{2}}|^{2})I\\ &+4\frac{c_{P}^{2}-2c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}\left(\nabla\log\sqrt{\rho_{1}}\otimes\nabla\log\sqrt{\rho_{1}}-\nabla\log\sqrt{\rho_{2}}\otimes\nabla\log\sqrt{\rho_{2}}\right)+c_{S}^{-2}(\nabla(\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}})\cdot\nabla c_{S}^{2})I\\ &+\left(2\frac{c_{P}^{2}-c_{S}^{2}}{c_{S}^{4}}\nabla c_{S}^{2}+c_{S}^{-2}\nabla c_{P}^{2}-\frac{12}{c_{P}^{2}-c_{S}^{2}}\nabla c_{S}^{2}\right)\otimes^{s}(\nabla\log\sqrt{\rho_{1}}-\nabla\log\sqrt{\rho_{2}})\\ =&0\quad\text{in }\Omega.\end{split}$
Applying divergence twice to the above equation, we obtain a fourth order linear differential equation of the form
$3\Delta^{2}(\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}})+\sum_{|\alpha|\leq 3}F_{\alpha}(\rho_{1},\rho_{2})D^{\alpha}(\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}})=0,\quad\text{in }\Omega.$
By the Unique Continuation Principle (cf. [17]) and the boundary determination results $\partial_{\nu}^{k}\rho_{1}=\partial_{\nu}^{k}\rho_{2}$, $k=0,1,2,3$ on $\partial\Omega$ (cf. [18]), we can conclude that
$\rho_{1}=\rho_{2}\quad\text{in }\overline{\Omega}.$
This completes the proof of the main theorem.

### v3

@@ Appendix A Proof of lemmas
$+\nabla\log\sqrt{\rho}\cdot\nabla c_{S}^{2}+c_{S}^{4}\nabla_{\beta}\log\sqrt{\rho}\left(2\frac{c_{P}^{2}-c_{S}^{2}}{c_{S}^{4}}\nabla_{\alpha}c_{S}^{2}+c_{S}^{-2}\nabla_{\alpha}c_{P}^{2}-\frac{12}{c_{P}^{2}-c_{S}^{2}}\nabla_{\alpha}c_{S}^{2}\right)\eta^{\alpha}\eta^{\beta}$
$+\text{terms that do not depend on $\rho$}$
$=$ $:c^{4}_{S}g^{\alpha\varepsilon}g^{\beta\delta}N_{\alpha\beta}\eta_{\varepsilon}\eta_{\delta}+C$
$=$ $g_{S}^{\alpha\varepsilon}g_{S}^{\beta\delta}N_{\alpha\beta}\eta_{\varepsilon}\eta_{\delta}+C,$
where $C$ does not depend on $\rho$. Recall that $g_{S}=c_{S}^{-2}g_{E}$, and consequently
$g_{S}^{\alpha\beta}=c_{S}^{2}g^{\alpha\beta}.$
One should be careful in the above calculation that $\eta^{\alpha}\partial_{x^{\alpha}}$ is tangent vector corresponding to the cotangent vector $\eta$ w.r.t. the Euclidean metric $g_{E}$. The corresponding tangent vector w.r.t. the metric $g_{S}$ should be $c_{S}^{2}\eta^{\alpha}\partial_{x^{\alpha}}$. Note that the second order tensor $N$ can be written in a coordinate-free way
@@
$\begin{split}&\mathrm{i}c_{S}^{-2}\left(2\mu\partial_{3}a_{1\alpha}+(\partial_{3}\mu)a_{1\alpha}-\mu g^{\beta\varepsilon}\frac{\partial g_{\alpha\varepsilon}}{\partial\tau}a_{1\beta}+\frac{1}{2}g^{\beta\varepsilon}\frac{\partial g_{\beta\varepsilon}}{\partial\tau}c_{S}^{-2}\mu a_{1\alpha}-c_{S}^{-1}\partial_{3}c_{S}\mu a_{1\alpha}\right)c_{S}^{2}\eta^{\alpha}\\ =&(\langle N,\eta\otimes\eta\rangle_{g_{S}}+C)J^{-1/2}c_{S}^{1/2}\rho^{1/2}.\end{split}$
Denote $D_{\tau}$ to be the covariant derivative along the geodesic. We consider the above equation along $\gamma$. If we denote
$a_{1i}=J^{-1/2}c_{S}^{1/2}\rho^{-1/2}\xi_{i},$
with a vector field $\xi=(\xi_{1},\xi_{2},\xi_{3})$ along $\gamma$, the above equation can be written as
@@
along $\gamma$. The proof is complete if one notices that $\langle\xi,\eta\rangle_{g_{S}}=B_{S}$. ∎
Acknowledgement. The author thanks the anonymous referees for very useful comments that helped improve this paper significantly.
Data Availability This article has no associated data.
Conflict of interest The authors have no conflict of interest to declare that are relevant to the content of this article.